OpenStudy (anonymous):

Jake can cut firewood in 6 fewer hours than skyler can. When they work together, it takes them 4 hours. How long would it take each of them to do the job alone?

OpenStudy (anonymous):

2x-6=4 2x=10 x=5 skyler x-6=5 x=11 jake x=5

OpenStudy (anonymous):

jake is x; skyler x-6

OpenStudy (anonymous):

x = Hours it takes Skyler to complete x - 6 = Hours it takes Jake to complete 1/(x-6) + 1/x = 1/4 Solve for x.

OpenStudy (anonymous):

Jake: 6 hours Skyler: 12 hours